# Extending Stable and Popular Matching Algorithms from Bipartite to Arbitrary Instances

### v1

@@
Abstract
@@ 1 Introduction
Preference based matching markets is an extensively studied topic in computer science, mathematics and economics literature. The extensive study of the area started back in 1962 after the seminal paper of Gale and Shapley [13], who introduced the stable marriage problem and showed that a stable matching always exists and can be found in linear time. Since then, countless applications and related models have been studied, see Manlove [27] for an overview. Applications include Resident allocation, University admission, Kidney exchange, job markets and much more.
@@ 1.1 Related Work
In the stable marriage problem, the goal is to match two classes of agents in a way such that there are no blocking edges, i.e. pairs of agents who mutually prefer each other to their partners. The stable marriage problem with ties and incomplete lists (smti) was first studied by Iwama et al.  [20], who showed the NP-hardness of Max-SMTI, which is the problem of finding a maximum size stable matching in an SMTI instance. Since then, several papers made progress to improve the approximability of the problem, e.g. Iwama et al. [21, 22] and Király [24]. Currently, the best ratio is $\frac{3}{2}$ by a polynomial time algorithm of McDermid [29], where the same ratio is attained by linear-time algorithms of Pauluch [31, 32] and Király [25, 26]. This $\frac{3}{2}$-approximation has been extended even to very general cases, including two-sided matroid constraints [9], $\gamma$-stability [7], where we are given $\gamma_{v}^{e}$ numbers for each edge $e$ and vertex $v$ such that these numbers describe how much improvement $v$ needs to consider $e$ a blocking edge, and critical relaxed stability [30], which is a generalization of Max-SMTI, where there is a set of critical agents, among whom we must match as much as possible and stability is relaxed in an appropriate way.
@@
As for the inapproximability of Max-SMTI, Halldórssson et al. [16] showed that it is NP-hard to approximate it within some constant factor. Later, inapproximability results have been improved, especially assuming stronger complexity theoretic conjectures. Yanagisawa [34] and Halldórsson et al. [17] showed that assuming the Unique Games Conjecture, there is no $\frac{4}{3}-\varepsilon$-approximation for any $\varepsilon>0$. By a recent work by Dudycz et al. [12] it follows that assuming the strong Unique Games Conjecture or the Small Set Expansion Hypothesis or the strong-UGC, there cannot even be a $\frac{3}{2}-\varepsilon$-approximation algorithm for Max-SMTI.
@@
The Stable Roommates problem with ties and incomplete preferences was studied by Irving and Manlove [19]. They showed that finding super stable matchings can be done in polynomial time. They posed as an open question whether the maximum size of a weakly stable matching can be approximated better than 2 or not. Of course, as deciding if a stable roommates instance with ties admits a weakly stable matching is NP-hard, the question makes sense in the following two senses. First of all, by a result of Tan [33], we know that half-integral weakly stable matchings (also called weakly stable partitions) always exist and can be found in polynomial time, even in the case of multigraphs [6]. Therefore, the question of whether we can approximate the maximum size of a weakly stable half-matching (or a weakly stable fractional-matching) better than 2 is still a very relevant open question. Secondly, as in [1], we can restrict the instances to ones without odd preference cycles (meaning in every odd cycle there is an agent that prefers its clockwise neighbour to its other neighbour, and there is another agent that prefers its anti-clockwise neighbour to its other neighbour), which implies that there always exists a weakly stable matching. Hence, for such instances, it makes sense to try to find a weakly stable (integral) matching, which approximates the size of the maximum size weakly stable matching better than 2. Here, a 2-approximation is trivial, as any weakly stable matching is a 2-approximation (and every weakly stable half-matching gives at least one weakly stable integral matching of the same size), which has already been observed by Irving and Manlove in their paper [19]. We show that both these problems admit a $\frac{3}{2}$-approximation algorithm, which is best possible under the strong Unique Games Conjecture, given the inapproximabilty results for the bipartite case.
@@
Popular matchings have been studied by many papers recently. It has been known since Gardenfors’ seminal paper in 1975 [14] that stable matchings are popular, and hence popular matchings always exist in bipartite instances. Kavitha [23] gave algorithms to find a maximum size popular matching, and a matching that is popular among the maximum size matchings. The ideas of her algorithms were similar to Király’s, that is, they used Gale-Shapley type proposal-rejection algorithms, where agents could have multiple levels. Finding a popular matching among the maximum weight matchings has also a similar algorithm as shown in Csáji et al. [8].
@@
Sadly, if the preferences of the agents may contain ties, a popular matching may not exist and finding a popular matching is NP-hard [11]. In the case of arbitrary graphs, i.e. stable roommates instances, finding a popular matching was shown to be NP-hard even with strict preferences by Gupta et al. [15].
@@
To cope with this hardness, Huang and Kavitha [18] studied popular mixed-matchings in the marriage and roommmates settings. They showed that a popular mixed-matching always exists and gave an algorithm to find a maximum weight popular mixed-matching in polynomial-time by using an LP-formulation for the problem. This notion of popular mixed-matchings were also studied by Brandt and Bullinger [5]. We introduce a new version of popularity for fractional-matchings, which we call popular fractional-matchings. We show that popular mixed-matchings may not be popular fractional-matchings, but our algorithm can find a popular fractional-matching that has maximum size among both the popular fractional- and popular mixed-matchings.
@@ 1.2 Our Contributions
We also extend the algorithms for finding maximum size popular matchings and popular matchings among the maximum weight matchings (in the case of strict preferences). Since popular matchings may not exist in the roommates model, and even deciding their existence is NP-hard, we propose a natural definition of popularity among fractional-matchings. This is somewhat similar to a previous notion of popular mixed-matchings [18], however, we show that a mixed-matching may not be popular with respect to our notion. We show that any stable half-matching is popular and that there always exists a maximum size popular fractional-matching that is half-integral. We also give an algorithm that finds such a matching. This happens to be also a popular mixed-matching which has maximum size among the popular mixed-matchings. Finally, we give an algorithm that finds a half-integral matching that is popular among the maximum weight fractional-matchings.
@@ 1.3 Paper Structure
We start by defining the problems we study in more detail in Section 2. Then, we describe the meta-algorithm in Section 3, which serves as a basis for all three of our algorithms. The approximation algorithm for Max-SRTI is described in Section 3.1, the algorithm for finding a maximum size popular fractional-matching in Section 3.2 and the algorithm for finding a popular half-integral matching among the maximum weight fractional-matchings in Section 3.3. Finally, we conclude in Section 4.
@@ 2.1 Weakly Stable and $\gamma$-stable Matchings
A fractional-matching $M$ is called weakly stable or just stable, if there is no blocking edge to $M$. The problem of finding a maximum size weakly stable matching in a bipartite graph with weak preferences, called Max-SMTI is a well studied problem that is NP-hard even to approximate within some constant, but admits a linear time $\frac{3}{2}$-approximation [26]. Here, we consider the problem of finding a maximum-size weakly stable fractional-matching in arbitrary graphs, which we call maximum size stable roommates problem with ties and incomplete preferences, Max-SRTI for short.
@@
Max-$\gamma$-fSRTI In: A graph $G=(V;E)$, $p_{v}()$ preference valuations for each $v\in V$, numbers $0<\gamma_{e}^{v}<\delta_{e}^{v}$ for each pair $(e,v)\in E\times V$ such that $v\in e$. Out: A maximum size $\gamma$-stable fractional-matching $M$
The special case of finding a maximum size weakly stable fractional-matching is called Max-SRTI.
@@ 2.2 Popular Matchings
For our popular matching related problems, we assume that the $p_{v}(e)$ values are different for each vertex $v\in V$, so they induce a strict preference order $\succ_{v}$ over its adjacent edges. Also, $e\succ_{v}\emptyset$ for any adjacent edge $e$ to $v$.
@@
For integral matchings, popularity is defined as follows. $M$ is popular, if $\Delta(M,N)=\sum_{v\in V}\mathrm{vote}_{v}(M(v),N(v))\geq 0$ for any matching $N$, where $M(v)$ and $N(v)$ denote the edge $v$ receives in $M$ or $N$ respectively, if there is any, otherwise they are defined to be $\emptyset$. Hence, a matching $M$ is popular, if it does not lose in a head-to-head comparison to any other matching.
We extend this definition in a natural way to fractional-matchings, which mirrors the definition of popularity for the case of many-to-many matching instances.
@@
For a feasible pairing $\phi$, we define $\mathrm{vote}_{v}(M,N,\phi_{v})=\sum_{(e,f)\in(E(v)\cup\{\emptyset\})^{2}}\phi_{v}(e,f)\cdot\mathrm{vote}_{v}(e,f)$. Finally, we let $\Delta(M,N,\phi)=\sum_{v\in V}\mathrm{vote}_{v}(M,N,\phi_{v})$ and $\Delta(M,N)=\min\big{\{}\Delta(M,N,\phi)\bigm{|}\text{$\phi$ is a feasible pairing}\big{\}}$.
@@
Max-PRI In: A graph $G=(V;E)$ with strict preference orders $\succ_{v}$ for each $v\in V$. Out: A maximum size popular fractional-matching $M$
@@ 2.3 Popular Mixed-Matchings and a common generalization
However, it is not a popular fractional-matching. Take $N$ with $N(u_{1},w_{1})=1,N(u_{2},w_{2})=N(u_{3},w_{2})=\frac{1}{2}$ and $N(e)=0$ otherwise. Then, the feasible pairings are unique for each agent and we have that $u_{1},w_{1}$ both have votes $-\frac{1}{2}$, $u_{2},w_{2}$ both have votes $+\frac{1}{2}$ and $u_{3}$ has vote $-\frac{1}{2}$ so $\Delta(M.N)<0$.
We relax the axioms of feasible pairings to allow comparisons that arise from the definition of popular mixed-matchings -- note that here an edge $e$ with $M(e)=N(e)=\frac{1}{2}$ may be paired with different edges, which is not allowed in a feasible pairing.
@@
    $\sum_{f\in E(v)\cup\{\emptyset\}}\phi_{v}(\emptyset,f)>0$Longrightarrow $\sum_{e\in E(v)}M(e)<1$,
@@
    $\sum_{f\in E(v)\cup\{\emptyset\}}\phi_{v}(f,\emptyset)>0$Longrightarrow $\sum_{e\in E(v)}N(e)<1$,
@@
Note that by changing the last axiom to $\phi_{u}(e,e,)=\phi_{v}(e,e)=\min\{M(e),N(e)\}$, changing 3 back to the stronger requirement $\sum_{f\in E(v)\cup\{\emptyset\}}\phi_{v}(\emptyset,f)=(\sum_{e\in E(v)}N(e)-M(e))^{+}$, and changing 4 to $\sum_{f\in E(v)\cup\{\emptyset\}}\phi_{v}(f,\emptyset)=(\sum_{e\in E(v)}M(e)-N(e))^{+}$, we get back the definition for a feasible pairing. Hence, any feasible pairing is sensible. For a pairing $\phi_{v}$ defined by $\phi_{v}(e,f)=M(e)N(f)$, $\phi_{v}(e,\emptyset)=M(e)(1-\sum_{f\in E(v)}N(f))$, $\phi_{v}(\emptyset,f)=(1-\sum_{e\in E(v)}M(e))N(f)$ for $e=(u,v)$, which is the pairing in the definition of popular mixed-matchings, we have that all axioms are clearly satisfied.
We say that a fractional-matching $M$ is extra-popular, if $\Delta(M,N,\phi)\geq 0$ for any sensible-pairing $\phi$.
@@ 2.4 Popular Maximum Weight Matchings and Popular Critical Matchings
Let us suppose that we have some weight function $\omega:E\to\mathbb{R}$ on the edges. By restricting the candidate fractional-matchings to only the maximum weight fractional-matchings, we can define a new version of popularity. We say that a maximum weight fractional-matching $M$ is a popular maximum weight fractional-matching, if we have $\Delta(M,N,\phi)\geq 0$ for any other maximum weight fractional-matching $N$ and feasible pairing $\phi$. We call the problem of finding such a matching Pop-MaxW-SRI.
Pop-MaxW-SRI In: A graph $G=(V;E)$ with strict preference orders $\succ_{v}$ for each $v\in V$ and a weight function $\omega:E\to\mathbb{R}$. Out: A popular maximum weight fractional-matching $M$.
In our algorithm, we reduce this problem to a slightly easier problem, where we consider critical fractional-matchings. To define these, let’s suppose we are given a set $\mathcal{C}\subseteq V$ of critical vertices. We say that a fractional-matching is critical, if it saturates every critical vertex, i.e. if $\sum_{e\in E(v)}M(e)=1$ for all $v\in\mathcal{C}$.
@@
Pop-Crit-SRI In: A graph $G=(V;E)$ with strict preference orders $\succ_{v}$ for each $v\in V$ and a set of critical vertices $\mathcal{C}\subseteq V$. Out: A popular critical fractional-matching $M$.

### v3

@@
Extending Stable and Popular Matching Algorithms from Bipartite to Arbitrary Instances
Gergely Csáji1,2
1Institute of Economics, Centre for Economic and Regional Studies (ELTE KRTK KTI), Hungary
2Department of Operations Research, Eötvös Loránd University (ELTE), Hungary
@@ 1 Introduction
Preference-based matching markets are a well-studied area in computer science, mathematics, and economics, with many practical applications. Research in this field began in 1962 with the influential work of Gale and Shapley, who introduced the stable marriage problem and showed that a stable matching always exists and can be found in linear time [15]. This discovery led to a large body of research on matching problems and their applications. For an overview, see Manlove [29]. Today, matching theory is used in many areas, including resident allocation, university admissions, kidney exchanges, job markets, and more, making it important for various decision-making tasks
@@ 1.1 Related Work
The stable marriage model - where the goal is to match two classes of agents in a way such that there are no blocking edges, i.e. pairs of agents who mutually prefer each other to their partners - was introduced by Gale and Shapley [15]. However, this work only considered complete and strict preferences. The problem was first extended for the case where ties and incomplete preferences are allowed by Iwama et al. [22], who showed the NP-hardness of Max-SMTI, which is the problem of finding a maximum size weakly stable matching in such an instance, where weakly stable refers to the fact that both agents need a strict improvement for a pair to block. Since then, several papers made progress to improve the approximability of the problem, e.g. Iwama et al. [23, 24] and Király [26]. Currently, the best ratio is $\frac{3}{2}$ by a polynomial time algorithm of McDermid [31], where the same ratio is attained by linear-time algorithms of Pauluch [33, 34] and Király [27, 28]. This $\frac{3}{2}$-approximation has been extended even to very general cases, including two-sided matroid constraints [10], $\gamma$-stability [8], where we are given $\gamma_{v}^{e}$ numbers for each edge $e$ and vertex $v$ such that these numbers describe how much improvement $v$ needs to consider $e$ a blocking edge, and critical relaxed stability [32], which is a generalization of Max-SMTI, where there is a set of critical agents, among whom we must match as much as possible and stability is relaxed in an appropriate way.
@@
As for the inapproximability of Max-SMTI, Halldórssson et al. [18] showed that it is NP-hard to approximate it within some constant factor. Yanagisawa [36] and Halldórsson et al. [19] showed that assuming the Unique Games Conjecture, there is no $\frac{4}{3}-\varepsilon$-approximation for any $\varepsilon>0$. In a recent work, Dudycz et al. [13] proved that assuming the strong Unique Games Conjecture or the Small Set Expansion Hypothesis, there cannot even be a $\frac{3}{2}-\varepsilon$-approximation algorithm for Max-SMTI.
@@
The Stable Roommates problem with ties and incomplete preferences was studied by Irving and Manlove [21]. They showed that finding super stable matchings (where even a weak improvement is enough for a pair to block) can be done in polynomial time. They posed as an open question whether the maximum size of a weakly stable matching can be approximated better than 2 or not. Of course, as deciding if a stable roommates instance with ties admits a weakly stable matching is NP-hard [21], the question can be interpreted in the following sense. By a result of Tan [35], we know that half-integral weakly stable matchings (also called weakly stable partitions) always exist and can be found in polynomial time, even in the case of multigraphs [7]. Therefore, the question of whether we can approximate the maximum size of a weakly stable half-matching (or a weakly stable fractional-matching) better than 2 is still a very relevant open question. We show that both these problems admit a $\frac{3}{2}$-approximation algorithm, which is best possible under the strong Unique Games Conjecture, given the inapproximabilty results for the bipartite case.
@@
Popular matchings have been studied by many papers recently. Intuitively, a matching is popular, if for any other matching, the number of agents who improve is at most the number of agents who get worse. It has been known since Gardenfors’ seminal paper in 1975 [16] that stable matchings are popular, and hence popular matchings always exist in bipartite instances. Kavitha [25] gave algorithms to find a maximum size popular matching, and a matching that is popular among the maximum size matchings. The ideas of her algorithms were similar to Király’s, that is, they used Gale-Shapley type proposal-rejection algorithms, where agents could have multiple levels. Finding a popular matching among the maximum weight matchings has also a similar algorithm as shown in Csáji et al. [9].
@@
Sadly, if the preferences of the agents may contain ties, a popular matching may not exist and finding a popular matching or reporting that none exist is NP-hard [12]. In the case of arbitrary graphs, i.e. stable roommates instances, finding a popular matching was shown to be NP-hard even with strict preferences by Gupta et al. [17] and Faenza at al. [14] independently.
@@
To cope with this computational obstacle, Huang and Kavitha [20] studied popular mixed-matchings in the marriage and roommmates settings. They showed that a popular mixed-matching always exists and gave an algorithm to find a maximum weight popular mixed-matching in polynomial-time by using an LP-formulation for the problem. This notion of popular mixed-matchings were also studied by Brandt and Bullinger [6]. We introduce a new version of popularity for fractional-matchings, which we call popular fractional-matchings. We show that popular mixed-matchings may not be popular fractional-matchings, but our algorithm can find a popular fractional-matching that has maximum size among both the popular fractional- and popular mixed-matchings.
@@ 1.2 Our Contributions
We also extend the algorithms for finding maximum size popular matchings and popular matchings among the maximum weight matchings (in the case of strict preferences). Since popular matchings may not exist in the roommates model, and even deciding their existence is NP-hard, we propose a natural definition of popularity among fractional-matchings. This is somewhat similar to a previous notion of popular mixed-matchings [20], however, we show that a mixed-matching may not be popular with respect to our notion. We show that there always exists a maximum size popular fractional-matching that is half-integral. We also give an algorithm that finds such a matching. This happens to be also a popular mixed-matching which has maximum size among the popular mixed-matchings. Finally, we give an algorithm that finds a half-integral matching that is popular among the maximum weight fractional-matchings.
@@ 1.3 Paper Structure
We start by defining the problems we study in more detail in Section 2. Then, we describe the meta-algorithm in Section 3, which serves as a basis for all three of our algorithms. The approximation algorithm for Max-fSRTI is described in Section 3.1, the algorithm for finding a maximum size popular fractional-matching in Section 3.2 and the algorithm for finding a popular half-integral matching among the maximum weight fractional-matchings in Section 3.3. Finally, we conclude in Section 4.
@@ 2.1 Weakly Stable and $\gamma$-stable Matchings
A fractional-matching $M$ is called weakly stable or just stable, if there is no blocking edge to $M$. The problem of finding a maximum size weakly stable matching in a bipartite graph with weak preferences, called Max-SMTI is a well studied problem that is NP-hard even to approximate within some constant, but admits a linear time $\frac{3}{2}$-approximation [28]. Here, we consider the problem of finding a maximum-size weakly stable fractional-matching in arbitrary graphs, which we call maximum size stable roommates problem with ties and incomplete preferences, or Max-fSRTI for short.
@@
Max-$\gamma$-fSRTI In: A graph $G=(V;E)$, $p_{v}:E(v)\to\mathbb{R}_{\geq 0}$ preference valuations for each $v\in V$, numbers $0<\gamma_{e}^{v}<\delta_{e}^{v}$ for each pair $(e,v)\in E\times V$ such that $v\in e$. Out: A maximum size $\gamma$-stable fractional-matching $M$
The special case of finding a maximum size weakly stable fractional-matching is called Max-fSRTI.
@@ 2.2 Popular Matchings
For our popular matching related problems, we assume that the $p_{v}(e)$ values are different for each vertex $v\in V$, so they induce a strict preference order $\succ_{v}$ over its adjacent edges. Also, we assume $e\succ_{v}\emptyset$ for any adjacent edge $e$ to $v$.
@@
For integral matchings, popularity is defined as follows. $M$ is popular, if
$\Delta(M,N)=\sum_{v\in V}\mathrm{vote}_{v}(M(v),N(v))\geq 0$
for any matching $N$, where $M(v)$ and $N(v)$ denote the edge $v$ receives in $M$ or $N$ respectively, if there is any, otherwise they are defined to be $\emptyset$. Hence, a matching $M$ is popular, if it does not lose in a head-to-head comparison to any other matching.
We extend this definition in a natural way to fractional-matchings, which mirrors the definition of popularity for the case of many-to-many matching instances [5].
@@
For a feasible pairing $\phi$, we define
$\mathrm{vote}_{v}(M,N,\phi_{v})=\sum_{(e,f)\in(E(v)\cup\{\emptyset\})^{2}}\phi_{v}(e,f)\cdot\mathrm{vote}_{v}(e,f).$
Finally, we let
$\Delta(M,N,\phi)=\sum_{v\in V}\mathrm{vote}_{v}(M,N,\phi_{v})$
and
$\Delta(M,N)=\min\big\{\Delta(M,N,\phi)\bigm|\text{$\phi$ is a feasible pairing}\big\}$
@@
Max-fPRI In: A graph $G=(V;E)$ with strict preference orders $\succ_{v}$ for each $v\in V$. Out: A maximum size popular fractional-matching $M$
@@ 2.3 Popular Mixed-Matchings and a Common Generalization
However, it is not a popular fractional-matching. Take $N$ with $N(u_{1},w_{1})=1,N(u_{2},w_{2})=N(u_{3},w_{2})=\frac{1}{2}$ and $N(e)=0$ otherwise. Then, the feasible pairings are unique for each agent and we have that $u_{1},w_{1}$ both have votes $\frac{1}{2}$, $u_{2},w_{2}$ both have votes $\frac{1}{2}$ and $u_{3}$ has vote $\frac{1}{2}$ so $\Delta(M.N)<0$.
We relax the axioms of feasible pairings to allow comparisons that arise from the definition of popular mixed-matchings - note that here an edge $e$ with $M(e)=N(e)=\frac{1}{2}$ may be paired with different edges, which is not allowed in a feasible pairing.
@@
    $\sum_{f\in E(v)\cup\{\emptyset\}}\phi_{v}(\emptyset,f)>0$ $\Longrightarrow$ $\sum_{e\in E(v)}M(e)<1$,
@@
    $\sum_{f\in E(v)\cup\{\emptyset\}}\phi_{v}(f,\emptyset)>0$ $\Longrightarrow$ $\sum_{e\in E(v)}N(e)<1$,
@@
Note that by changing the last axiom to $\phi_{u}(e,e,)=\phi_{v}(e,e)=\min\{M(e),N(e)\}$, changing 3. back to the stronger requirement $\sum_{f\in E(v)\cup\{\emptyset\}}\phi_{v}(\emptyset,f)=(\sum_{e\in E(v)}N(e)-M(e))^{+}$, and changing 4. to $\sum_{f\in E(v)\cup\{\emptyset\}}\phi_{v}(f,\emptyset)=(\sum_{e\in E(v)}M(e)-N(e))^{+}$, we get back the definition for a feasible pairing. Hence, any feasible pairing is sensible. For a pairing $\phi_{v}$ defined by $\phi_{v}(e,f)=M(e)N(f)$, $\phi_{v}(e,\emptyset)=M(e)(1-\sum_{f\in E(v)}N(f))$, $\phi_{v}(\emptyset,f)=(1-\sum_{e\in E(v)}M(e))N(f)$ for $e=(u,v)$, which is the pairing in the definition of popular mixed-matchings, we have that all axioms are clearly satisfied.
We say that a fractional-matching $M$ is extra-popular, if $\Delta(M,N,\phi)\geq 0$ for any fractional matching $N$ and any sensible-pairing $\phi$ between $M$ and $N$.
@@ 2.4 Popular Maximum Weight Matchings and Popular Critical Matchings
Let us suppose that we have some weight function $\omega:E\to\mathbb{R}$ on the edges. By restricting the candidate fractional-matchings to only the maximum weight fractional-matchings, we can define a new version of popularity. We say that a maximum weight fractional-matching $M$ is a popular maximum weight fractional-matching, if we have $\Delta(M,N,\phi)\geq 0$ for any other maximum weight fractional-matching $N$ and feasible pairing $\phi$. We call the problem of finding such a matching MaxW-fPRI.
MaxW-fPRI In: A graph $G=(V;E)$ with strict preference orders $\succ_{v}$ for each $v\in V$ and a weight function $\omega:E\to\mathbb{R}$. Out: A popular maximum weight fractional-matching $M$.
In our algorithm, we reduce this problem to a slightly easier problem, where we consider critical fractional-matchings. To define these, let us suppose we are given a set $\mathcal{C}\subseteq V$ of critical vertices. We say that a fractional-matching is critical, if it saturates every critical vertex, i.e. if $\sum_{e\in E(v)}M(e)=1$ for all $v\in\mathcal{C}$.
@@
Crit-fPRI In: A graph $G=(V;E)$ with strict preference orders $\succ_{v}$ for each $v\in V$ and a set of critical vertices $\mathcal{C}\subseteq V$. Out: A popular critical fractional-matching $M$.